OpenStudy (anonymous):

What is the value of w? 2(2w-6)=6(w-6)+14 w=?

OpenStudy (anonymous):

2(2w-6)=6(w-6) +14 4w-12=6w-36+14 4w=6w-10 -2w-10 w=5